# On Early Detection of Hallucinations in Factual Question Answering

### v2

@@ 3.2. Artifacts for Detecting Hallucinations
We consider four sets of model artifacts for detecting hallucinations.
@@ 3.2.1. Softmax probabilities
We posit that the distribution of Softmax probabilities can be used to detect hallucinations. Specifically, we hypothesize that the Softmax distribution has a higher entropy when the model is hallucinating. A higher entropy in the Softmax distribution means that the model is “less sure” about its prediction. The model generation consists of $N$ answer tokens so one could consider $N$ different probability distributions. For simplicity, we mainly focus on the distribution at the first generated token, but analyze the effect of focusing on other tokens in Appendix A.
Figure 1(a) shows an example of difference in Softmax probabilities. The figure is generated by passing two questions from the TriviaQA (joshi-etal-2017-triviaqa, 15) dataset through the FAL-40B model. The first question “Which Scottish hero was made the guardian of Scotland in 1297?” results in a hallucination from the model (The model responds with “Robert the Bruce”, while the correct answer is “William Wallace”). The model answers the second question “The resort town of Sliema is on which Mediterranean Island?” correctly (Malta). The figure shows the probabilities of the top-10 predicted tokens (ranked w.r.t. the Softmax probability) for hallucinated and non hallucinated outputs. The distribution of the probabilities are notably different.
(kadavath2022language, 16) (kadavath2022language, 16) also test a similar hypothesis that the entropy in the generated tokens could be a signal of hallucinations. However, their approach is slightly different from ours. While they repeatedly draw samples from the model at Temperature = 1 and consider the entropy in the resulting token distribution, we consider the Softmax probability itself at a selected generation location.
@@ 3.2.2. Feature attributions
Feature attributions, that is, how important a feature was towards a certain prediction are often used to inspect the behavior of the model and find potential problematic patterns (doshi2017towards, 7, 21). Building on these insights, we posit that when answering the questions correctly, the model would focus on few input tokens. For example, when answering “Berlin” to “What is the capital of Germany?” the model would focus on “Germany”. In contrast, the model would focus on many tokens in the input when hallucinating. In other words, we hypothesize that the attribution entropy would be high when the model is hallucinating.
@@
There is a plethora of methods for generating $\boldsymbol{\Phi}_{i}$  (gilpin2018explaining, 11, 13). In this work, we use the Integrated Gradients (IG) (sundararajan2017axiomatic, 31). We select IG for the following reasons:
  1. (1)
    It provides attractive theoretical properties, e.g., efficiency meaning that the sum of all feature attributions equals the output Softmax probability.
  2. (2)
    By leveraging gradients, it runs faster than related methods like Kernel SHAP (lundberg2017unified, 22).
  3. (3)
    Unlike methods like Layerwise Relevance Propagation (montavon2019layer, 24) that require architecture-specific implementations, IG can operate on any architecture as long as the model gradients are available.
Using the same inputs from Section 3.2.1, we show the IG input token attributions for the hallucinated and non-hallucinated generations in Figure 1(b). The figure shows a clear difference in distributions: For the non-hallucinated output, the LLM focuses on key tokens in the input (“Mediterranean” and “town”). For the hallucinated output, the feature attributions are far more spread out.
@@ 3.2.3. Self-attention and Hidden Activations
Formally, given a Transformer language model with $L$ layers, let $\mathbf{S}_{\ell}$ denote the self-attention at layer $\ell\in{1,\ldots,L}$ and $\mathbf{H}_{\ell}$ denote the fully-connected activations. We focus specifically on $\mathbf{S}_{\ell}^{(q_{M},a_{1})}$ and $\mathbf{H}_{\ell}^{(q_{M},a_{1})}$ which represented the self-attention and fully-connected activations between the final token of the input question $\mathbf{Q}$ and the first token of the response $a_{1}$. Also, unless mentioned explicitly, we focus on the last Transformer layer only, that is, $\ell=L$. This choice was made based on the preliminary experiments that showed the last layer to provide the most promising performance (see Section 5.2).
@@ 3.3. Hallucination Detection Classifiers
Since our goal is to assign a hallucination / non-hallucination label to the model generations, in this section, we describe how to use the artifacts detailed in Section 3.2 to arrive at these labels.
@@
Classifier architecture. The classifiers using IG attributions consist of a 4 layer Gated Recurrent Unit network with $25\%$ dropout at each layer. We use a recurrent network instead of feed-forward because the dimensionality of attributions is different for each input (one attribution score per input token), and instead of a Transformer because of the relatively small amount of training data. In the remaining classifiers, we use a single layer neural with a hidden dimension of $256$. For each dataset, we train and evaluate on a random 80/20 split.
@@ 4. Experimental Setup
In this section, we describe the datasets, LLMs and the configurations used in our experiments. We also report the accuracy of LLMs in carrying out the base task (answering questions without hallucinations).
@@ 4.1. Datasets
We use two QA datasets: the T-REx dataset from (elsahar-etal-2018-rex, 8) and the TriviaQA dataset from (joshi-etal-2017-triviaqa, 15).
@@ 4.1.1. T-REx
The T-REx dataset consists of relationship triplets consisting of pairs of entities and their relationships, e.g., (France, Paris, Capital of) and (Tsar Peter I, Moscow, Born in). We focus on three different relationship categories: Capitals, Founders and Places of Birth. In our experiments, we report results separately for each of the three categories since we found different patterns across the categories.
@@ 4.1.2. TriviaQA
TriviaQA is a reading comprehension dataset consists of a set of $650,000$ trivia question, answer and evidence tuples. Evidence documents contain supporting information about the answer. We only use the closed book setting (liang2022holistic, 19) where the model is only provided with questions without any supporting information. Some example questions from the dataset are:
@@
(b) Softmax probability
(c) Self-attention
@@
Figure 3. [FAL-40B on TriviaQA dataset] Differences in the model artifacts for hallucinated v.s. non-hallucinated outputs based on (left to right) IG attributions, Softmax probabilities, self-attention and activations of the fully-connected layer. The top row shows the differences using the 2-dimensional TSNE projections whereas the bottom row shows the cumulative distribution function of the entropy.
@@
TriviaQA 0.599 0.568 0.480
Capitals 0.410 0.676 0.434
Founders 0.568 0.438 0.446
Birth Place 0.483 0.451 0.515
Combined 0.433 0.570 0.491
@@
TriviaQA 0.715 0.633 0.605
Capitals 0.683 0.697 0.637
Founders 0.615 0.677 0.668
Birth Place 0.666 0.620 0.664
Combined 0.698 0.676 0.621
@@
TriviaQA 0.711 0.659 0.719
Capitals 0.720 0.726 0.720
Founders 0.734 0.686 0.711
Birth Place 0.812 0.612 0.815
Combined 0.784 0.712 0.790
@@
TriviaQA 0.722 0.644 0.722
Capitals 0.736 0.716 0.707
Founders 0.714 0.725 0.731
Birth Place 0.800 0.777 0.767
Combined 0.799 0.738 0.821
@@
Table 2. [Larger model variants] Test AUROC of binary classifiers in detecting hallucinations using (a) IG attributions, (b) Softmax probability distribution, (c), self-attention and (4) fully-connected activations.
@@
TriviaQA 0.621 0.547 0.549
Capitals 0.355 0.341 0.409
Founders 0.409 0.483 0.548
Birth Place 0.425 0.445 0.430
Combined 0.409 0.422 0.536
@@
TriviaQA 0.680 0.655 0.646
Capitals 0.650 0.672 0.684
Founders 0.660 0.661 0.726
Birth Place 0.591 0.650 0.646
Combined 0.673 0.671 0.728
@@
TriviaQA 0.710 0.668 0.663
Capitals 0.763 0.738 0.686
Founders 0.700 0.720 0.660
Birth Place 0.768 0.647 0.723
Combined 0.758 0.756 0.717
@@
TriviaQA 0.740 0.707 0.703
Captials 0.748 0.754 0.739
Founders 0.720 0.719 0.703
Birth Place 0.711 0.716 0.789
Combined 0.770 0.772 0.719
@@
Table 3. [Smaller model variants] Test AUROC of binary classifiers in detecting hallucinations using (a) IG attributions, (b) Softmax probability distribution, (c), self-attentinon and (4) fully-connected activations.
4.2. Models
We analyze responses from three different models: OpenLLaMA (LAM), [^1] https://huggingface.co/docs/transformers/main/model_doc/open-llama OPT [^2] https://huggingface.co/docs/transformers/v4.19.2/en/model_doc/opt and Falcon (FAL). [^3] https://huggingface.co/docs/transformers/main/model_doc/falcon All three models come with different number of parameters and we consider two different sizes for each model: LAM-7B (7 billion parameters) and LAM-13B; OPT13B and OPT30B; FAL-7B and FAL-40B. We consider different variants to study the effect of model size on the hallucination detection performance.
4.3. Infrastructure and Parameters
All experiments were ran on an Amazon SageMaker ml.g5.12xlarge instance with 192 GB RAM and 4x24GB NVIDIA A10G GPUs. Self-attentions and fully-connected activations were captured during the inference forward pass using the Amazon SageMaker Debugger (rauschmayr2021debugger, 29) service.
After passing in the prompt, generations are performed by sampling the most likely token according to the Softmax probability. This corresponds to using a Temperature = 0. We continue generating until an <end of text> tokens is generated or the generation is $100$ tokens long.
We compute the Integrated Gradients attribution using the Captum library (kokhlikyan2020captum, 17). The IG parameters are configured as follows. We use a baseline of all 00s and the number of IG iterations is set of $50$. Given an output token, the corresponding IG attributions for each input token are a vector of the same dimensionality as the input token embeddings. We convert the vector scores to the a per token scalar score by using the L2 norm reduction, which has been shown to provide similar or better performance to other reduction strategies (zafar2021lack, 35).
Hallucination classifiers are trained with a batch size of $128$ for $1,000$ iterations. We use Adam optimizer with a learning rate of $0.0001$ and weight decay of $0.01$.
4.4. Question Answering Accuracy
Before moving on to hallucination detection, we first analyze the performance of the models in correctly answering the questions i.e., how often the models hallucinate. Table 1 shows the accuracy of different models. The models showed a range of performance across each task. All models performed best on the TriviaQA tasks, and worst on the Birth Place task.
Surprisingly, larger models did not always perform better. On the subject specific tasks from T-REx, smaller models often performed as well or better than their larger counterparts (6 out of 9 times). On the more general TriviaQA task, larger models consistently performed better. On all tasks, FAL-40B significantly outperformed all other models.
Further, while larger models on average performed better at general knowledge tasks, variation in performance is more strongly correlated with model type rather than size. For example, while LAM-13B outperformed its smaller variants LAM-7B, and similarly OPT-30B outperformed its smaller variant, both LAM models outperformed both OPT models.
@@ 5. Results
In this section we first qualitatively investigate the differences in distributions of hallucinations and non-hallucinations based on the four artifacts that we identified in Section 3.2. We then quantitatively evaluate the performance of hallucination classifiers trained on these artifacts.
@@ 5.1. Qualitative analysis
Recall our hypotheses from Section 3.2 that hallucinated v.s. non-hallucinated outputs have a different distribution of Softmax probabilities, IG attributions, self attention and fully-connected activations. We also posited that when hallucinating, the models would be “unsure” about which tokens to generate next and would have a rather spread out input attribution, resulting in higher entropy for Softmax probabilities and IG token attributions, respectively.
Examining FAL-40B on the TriviaQA dataset (Figure 3), we find that the hallucinated outputs indeed show higher entropy than non-hallucinated ones. The trend is much smaller (but visually discernible) for IG attributions and fully-connected activations, and non-existent for self attention. Further, by projecting each set of artifacts into two dimensions using TSNE, we see distinctive clusterings of hallucinated results for all but the integrated gradients. The model internal states, attention and fully-connected activations, show distinct clusters of hallucinated outputs.
@@
The results on other datasets and models Appendix A are mixed. The distributions of input token entropy shows a stronger correlation to dataset type than to model. For topic specific datasets, entropy is on average lower for hallucinated results, while it is slightly higher for hallucinated results on TriviaQA. The entropies of the Softmax outputs are less stable across both datasets and models. While most cases show slightly higher entropy for hallucinated results, the difference is inconsistent. Further, in both the input token and output Softmax cases, the differences, while present, are relatively small. We find similarly small difference in entropy along all generated tokens (see Figure 4). 2-dimensional TSNE plots also show similarly mixed trends.
While mixed, the results still show that in many cases, the reduced distributions of model artifacts (either to 2 dimensions via TSNE or to 1 dimension via entropy) show discernible differences between hallucinated and non-hallucinated outputs. Next we investigate if we can train accurate hallucination detectors by using these artifacts in their original form i.e., without reductions.

### v3

@@ 3.2. Artifacts for Detecting Hallucinations
We consider four model artifacts for detecting hallucinations.
@@ 3.2.1. Softmax probabilities
We posit that the Softmax probability distribution can be used to detect hallucinations. Specifically, following the analysis in (kadavath2022language), we hypothesize that the Softmax distribution has a higher entropy when the model is hallucinating. A higher entropy means that the model is “less sure” about its prediction. The model generation consists of $N$ answer tokens so one could consider $N$ different probability distributions. For simplicity, we mainly focus on the distribution at the first generated token, but analyze the effect of focusing on other tokens in §5.2.
Figure 2(a) shows an example of difference in Softmax probabilities. The figure is generated by passing two questions from the TriviaQA (joshi-etal-2017-triviaqa) dataset through the Falcon-40B model. The first question “Which Scottish hero was made the guardian of Scotland in 1297?” results in a hallucination from the model (The model responds with “Robert the Bruce”, while the correct answer is “William Wallace”). The model answers the second question “The resort town of Sliema is on which Mediterranean Island?” correctly (Malta). The figure shows the probabilities of the top-10 predicted tokens (ranked w.r.t. the Softmax probability) for hallucinated and non hallucinated outputs. The distribution of the probabilities are notably different.
kadavath2022language (kadavath2022language) also test a similar hypothesis that the entropy in the generated tokens could be a signal of hallucinations. However, their approach is slightly different from ours. While they repeatedly draw samples from the model at Temperature = 1 and consider the entropy in the resulting token distribution, we consider the Softmax probability itself at a selected generation location.
@@ 3.2.2. Feature attributions
Feature attributions, that is, how important a feature was towards a certain prediction are often used to inspect the behavior of the model and find potential problematic patterns (doshi2017towards; lipton2018mythos). Building on these insights, we posit that when answering the questions correctly, the model would focus on few input tokens. For example, when answering “Berlin” to “What is the capital of Germany?” the model would focus on “Germany”. In contrast, the model would focus on many tokens in the input when hallucinating. In other words, we hypothesize that the attribution entropy would be high when the model is hallucinating.
@@
There is a plethora of methods for generating $\boldsymbol{\Phi}_{i}$  (gilpin2018explaining; guidotti2018survey). In this work, we use the Integrated Gradients (IG) (sundararajan2017axiomatic). We select IG for the following reasons: It provides attractive theoretical properties, e.g., efficiency meaning that the sum of all feature attributions equals the output Softmax probability. By leveraging gradients, it runs faster than related methods like Kernel SHAP (lundberg2017unified). Unlike methods like Layerwise Relevance Propagation (montavon2019layer) that require architecture-specific implementations, IG can operate on any architecture as long as the model gradients are available.
We show the IG input token attributions for a hallucinated and non-hallucinated generations in Figure 2(b). The figure shows a clear difference in distributions: For the non-hallucinated output, the LLM focuses on key tokens in the input (“town” and “island”). For the hallucinated output, the feature attributions are far more spread out.
@@ 3.2.3. Self-attention and Hidden Activations
Formally, given a Transformer language model with $L$ layers, let $\mathbf{S}_{\ell}$ denote the self-attention at layer $\ell\in{1,\ldots,L}$ and $\mathbf{H}_{\ell}$ denote the fully-connected activations. We focus specifically on $\mathbf{S}_{\ell}^{(q_{M},a_{1})}$ and $\mathbf{H}_{\ell}^{(q_{M},a_{1})}$ which represent the self-attention and fully-connected activations between the final token of the input question $\mathbf{Q}$ and the first token of the response $a_{1}$. Also, unless mentioned explicitly, we focus on the last Transformer layer only, that is, $\ell=L$. This choice was made based on the preliminary experiments that showed the last layer to provide the most promising performance (see §5.2).
@@ 3.3. Hallucination Detection Classifiers
Since our goal is to assign a hallucination / non-hallucination label to the model generations, we now describe how to use the artifacts detailed in §3.2 to arrive at this binary label.
@@
Classifier architecture. The classifiers using IG attributions consist of a 4 layer Gated Recurrent Unit network with $25\%$ dropout at each layer. We use a recurrent network instead of feed-forward because the dimensionality of attributions is different for each input (one attribution score for each token in the question), and instead of a Transformer because of the relatively small amount of training data. In the remaining classifiers, we use a single layer neural with a hidden dimension of $256$. For each dataset, we train and evaluate on a random 80/20 split.
@@ 4. Experimental Setup
In this section, we describe the datasets, LLMs and the configurations used in our experiments. We also report the accuracy of LLMs in carrying out the base task (answering questions accurately).
@@ 4.1. Datasets
We use the following two QA datasets: the T-REx dataset (elsahar-etal-2018-rex) the TriviaQA dataset (joshi-etal-2017-triviaqa).
@@ 4.1.1. T-REx
The T-REx dataset consists of relationship triplets containing pairs of entities and their relationships, e.g., (France, Paris, Capital of) and (Tsar Peter I, Moscow, Born in). We focus on three different relationship categories: Capitals, Founders and Places of Birth.
@@ 4.1.2. TriviaQA
TriviaQA is a reading comprehension dataset consists of a set of $650,000$ trivia question, answer and evidence tuples. Evidence documents contain supporting information about the answer. We only use the closed book setting (liang2022holistic) where the model is only provided with questions without any supporting information. Some example questions from the dataset are:
@@
(b) Softmax probabilities
(c) Self-attention scores
@@
Figure 3. [FAL-40B on TriviaQA dataset] 2D TSNE projections of the model artifacts for hallucinations and non-hallucinations.
4.2. Models
We analyze responses from three different models: OpenLLaMA (LAM), [^1] https://huggingface.co/docs/transformers/main/model_doc/open-llama OPT (OPT) [^2] https://huggingface.co/docs/transformers/v4.19.2/en/model_doc/opt and Falcon (FAL). [^3] https://huggingface.co/docs/transformers/main/model_doc/falcon All three models come with different size-based variants. We consider two different sizes for each model: LAM-7B (7 billion parameters) and LAM-13B; OPT-6.7B and OPT-30B; FAL-7B and FAL-40B. We consider different variants to study the effect of model size on the hallucination detection performance.
4.3. Infrastructure and Parameters
All experiments were ran on an Amazon SageMaker ml.g5.12xlarge instance with $192$ GB RAM and 4x NVIDIA A10G GPUs. Self-attention and fully-connected activations were captured using Amazon SageMaker Debugger (rauschmayr2021debugger).
Generations are performed by sampling the most likely token according to the Softmax probability. This corresponds to using a Temperature = 0. We continue generating until an <end of text> tokens is generated or the generation is $100$ tokens long.
We compute the Integrated Gradients attribution using Captum (kokhlikyan2020captum). We use a baseline of all 00s and the number of IG iterations is set of $50$. Given an output token, the corresponding IG attributions for each input token are a vector of the same dimensionality as the input token embeddings. We convert the vector scores to the a per token scalar score by using the L2 norm reduction, which has been shown to provide similar or better performance to other reduction strategies (zafar2021lack).
Hallucination classifiers are trained with a batch size of $128$ for $1,000$ iterations. We use Adam optimizer with a learning rate of $10^{-4}$ and weight decay of $10^{-2}$.
4.4. Question Answering Accuracy
Before moving on to hallucination detection, we first analyze the performance of the models in correctly answering the questions i.e., how often the models hallucinate. Table 1 shows the accuracy of different models. The models showed a range of performance across each task. All models performed best on the TriviaQA tasks, and worst on the Birth Place task. Surprisingly, larger models did not always perform better. On the subject specific tasks from T-REx, smaller models often performed as well or better than their larger counterparts (6 out of 9 times). On the more general TriviaQA task, larger models consistently performed better. On all tasks, FAL-40B significantly outperformed all other models. Further, while larger models on average performed better at general knowledge tasks, variation in performance is more strongly correlated with model type rather than size. For example, while LAM-13B outperformed its smaller variants LAM-7B, and similarly OPT-30B outperformed its smaller variant, both LAM models outperformed both OPT models.
4.5. Accuracy of the hallucination heuristic
Recall from §3.1 that we consider a model answer $\mathbf{A}$ to be a hallucination if $\mathbf{R}\not\subseteq\mathbf{A}$. To evaluate this heuristic, we randomly sample $100$ TriviaQA generations with FAL-40B. For each generation, each of the three authors independently labeled it as (non-) hallucination. Based on the majority vote, the heuristic was correct in $98/100$ cases. The cases where the heuristic was incorrect are:
Question: In Greek mythology, one of the 12 Labours of Hercules was to produce ‘what’ item belonging to Amazonian queen Hippolyte?
Ref. Answer: Magical Girdle
Model Answer: The girdle of Hippolyte was a girdle that was worn by the Amazonian queen Hippolyte. It was one of the 12 Labours of Hercules.
Heuristic label: Hallucination
Human label: Not a hallucination
Question: Which British Prime Minister was the 1st Earl of Stockton?
Ref. Answer: Harold Macmillan or Earl of Stockton
Model Answer: Which British Prime Minister was the 1st Earl of Stockton?
Heuristic label: Not a hallucination
Human label: Hallucination
@@ 4.5. Accuracy of the hallucination heuristic
TriviaQA 0.60 0.57 0.48
Capitals 0.41 0.68 0.43
Founders 0.57 0.44 0.45
Birth Place 0.48 0.45 0.51
Combined 0.43 0.57 0.49
@@
TriviaQA 0.71 0.63 0.60
Capitals 0.68 0.69 0.63
Founders 0.61 0.67 0.66
Birth Place 0.66 0.62 0.66
Combined 0.69 0.67 0.62
@@
TriviaQA 0.71 0.65 0.71
Capitals 0.72 0.72 0.72
Founders 0.73 0.68 0.71
Birth Place 0.81 0.61 0.81
Combined 0.78 0.71 0.79
@@
TriviaQA 0.72 0.64 0.72
Capitals 0.73 0.71 0.70
Founders 0.71 0.72 0.73
Birth Place 0.80 0.77 0.76
Combined 0.79 0.73 0.82
@@
Table 2. [Larger model variants] Test AUROC of binary classifiers in detecting hallucinations.
@@
TriviaQA 0.62 0.54 0.54
Capitals 0.35 0.34 0.40
Founders 0.40 0.48 0.54
Birth Place 0.42 0.44 0.43
Combined 0.40 0.42 0.53
@@
TriviaQA 0.68 0.65 0.64
Capitals 0.65 0.67 0.68
Founders 0.66 0.66 0.72
Birth Place 0.59 0.65 0.64
Combined 0.67 0.67 0.72
@@
TriviaQA 0.71 0.66 0.66
Capitals 0.76 0.73 0.68
Founders 0.70 0.72 0.66
Birth Place 0.76 0.64 0.72
Combined 0.75 0.75 0.71
@@
TriviaQA 0.74 0.70 0.70
Captials 0.74 0.75 0.73
Founders 0.72 0.71 0.70
Birth Place 0.71 0.71 0.78
Combined 0.77 0.77 0.71
@@
Table 3. [Smaller model variants] Test AUROC of binary classifiers in detecting hallucinations.
@@ 5. Results
We first qualitatively investigate the differences in distributions of hallucinations and non-hallucinations and then quantitatively evaluate the performance of hallucination classifiers.
@@ 5.1. Qualitative analysis
Recall our hypotheses that hallucinating and non-hallucinating generations differ have different distributions of generation artifacts, namely, Softmax probabilities, IG attributions, self attention and fully-connected activations (§3.2). Examining FAL-40B on the TriviaQA dataset (Figure 3), we find that hallucinated outputs indeed tend to be distributed differently from non-hallucinated outputs when considering Softmax probabilities, self-attention scores and fully-connected activations. The IG attributions however do not show much difference in distributions.
@@
Figure 5. [FAL-40B on TriviaQA dataset] AUROC of the hallucination detectors using self-attention and fully-connected activations at different layers. The performance is better at later layers but has diminishing returns.
The results when comparing the entropy of distributions, and other datasets and models are mixed (figures omitted due to lack of space). The distributions of input token entropy showed a stronger correlation to dataset type than to model. For topic specific datasets, entropy is on average lower for hallucinated results, while it is slightly higher for hallucinated results on TriviaQA. The entropies of the Softmax outputs are less stable across both datasets and models. While most cases show slightly higher entropy for hallucinated results, the difference is inconsistent. Further, in both the input token and output Softmax cases, the differences, while present, are relatively small. We find similarly small difference in entropy along all generated tokens (see Figure 4). 2-dimensional TSNE plots also show similarly mixed trends.
While mixed, the results show that in many cases, the reduced distributions of model artifacts (either to 2 dimensions via TSNE or to 1 dimension via entropy) show visually discernible differences between hallucinated and non-hallucinated outputs. Next we investigate if we can train accurate hallucination detectors by using these artifacts in their original form i.e., without reductions.